# Context parroting: A simple but tough-to-beat baseline for foundation models in scientific machine learning

### v1

@@ Abstract
Recently-developed time series foundation models for scientific machine learning exhibit emergent abilities to predict physical systems. These abilities include zero-shot forecasting, in which a model forecasts future states of a system given only a short trajectory as context. Here, we show that foundation models applied to physical systems can give accurate predictions, but that they fail to develop meaningful representations of the underlying physics. Instead, foundation models often forecast by context parroting, a simple zero-shot forecasting strategy that copies directly from the context. As a result, a naive direct context parroting model scores higher than state-of-the-art time-series foundation models on predicting a diverse range of dynamical systems, at a tiny fraction of the computational cost. We draw a parallel between context parroting and induction heads, which explains why large language models trained on text can be repurposed for time series forecasting. Our dynamical systems perspective also ties the scaling between forecast accuracy and context length to the fractal dimension of the attractor, providing insight into the previously observed in-context neural scaling laws. Context parroting thus serves as a simple but tough-to-beat baseline for future time-series foundation models and can help identify in-context learning strategies beyond parroting.
@@ 1 Introduction
A key test of generalization in scientific machine learning (SciML) is the ability to forecast future states of a new physical system. Prior SciML approaches primarily focused on developing specialized forecasting models trained extensively on specific dynamical systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. However, the generality of these models is limited by the amount of system-specific data available, motivating the recent development of time-series foundation models [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], which are trained on vast amounts of observed and simulated time series, and which can subsequently perform zero-shot forecasts for any time series—including those generated by previously-unseen dynamical systems. Interestingly, it was recently found that, when available historical data is limited, time-series foundation models outperform classical deep learning models in forecasting chaotic dynamical systems [42].
What mechanisms do time-series foundation models use to make zero-shot forecasts, and why they are effective for dynamical systems? It was recently observed that one such foundation model, Chronos [35], often employs an extremely simple strategy when forecasting chaotic systems [42]. The strategy, context parroting, scans the context for repeating motifs and copies the part of the context following the best-matching motif as its prediction (Fig. 1). This can be viewed as a kind of “in-context nearest neighbor” algorithm, which has zero trainable parameters and thus requires no training. So how good is context parroting as a zero-shot forecasting strategy on its own?
Here, we compare context parroting with three state-of-the-art time series foundation models: Chronos [35], TimesFM [37], and TimeMoE [39] on the challenging task of forecasting chaotic systems. We find that context parroting outperforms all three foundation models in both zero-shot forecast accuracy and inference cost, especially for longer context windows. Our results suggest that current time series foundation models fail to extract physical insight from dynamical systems, and that their apparent performance is largely due to benchmark tasks that may be solved by simple repetition.
@@
    Introduce context parroting as a simple but tough-to-beat baseline for zero-shot forecasting of dynamical systems
@@
    Show that context parroting outperforms time-series foundation models in predicting chaotic systems
@@
    Explain the in-context neural scaling law between forecast accuracy and context length, linking the scaling coefficient to the fractal dimension of the underlying attractor
@@ 2.1 Foundation models for science
Foundation models have recently been introduced for many scientific machine-learning tasks [43], including partial differential equations [44, 45, 46, 47, 48, 49, 50], neuroscience [51, 52, 53], and weather forecasting [54, 55]. However, most of these foundation models remain a black box and haven’t provided interpretable insights into the underlying process. Here, we analyze context parroting as a simple mechanism coming out of time-series foundation models, noting its strengths and weaknesses as a zero-shot forecasting strategy. This strategy, and the insights gained here, can potentially be applied to other scientific tasks.
2.2 In-context neural scaling law
Neural scaling laws describe the relation between the performance of a neural network and certain resources, such as model size, data size, or the amount of compute [56, 57, 58, 59]. Such scaling laws allow practitioners to predict the performance of yet-to-be-trained models based on the available resources and allocate them optimally [60].
@@ 2.2 In-context neural scaling law
When applying LLMs to forecast dynamical systems, Liu et al. [61] recently observed an in-context neural scaling law, in which the test loss decreases with the context window following a power law. Here, we show that this in-context neural scaling law can be reproduced when using context parroting to predict dynamical systems, and the scaling coefficient can be linked to an invariant property of the underlying dynamic process (the fractal dimension of the chaotic attractor). This finding shows that neural scaling laws are intrinsically linked to invariant properties of the process generating the data, and the theory can potentially be generalized to other models and tasks (e.g., can we estimate the “fractal dimension” of language from the neural scaling law of LLMs?).
2.3 In-context learning and induction heads
@@
Induction heads are computational circuits that naturally emerge in simple transformers through training, and they have been hypothesized to underlie a lot of the in-context learning ability of foundation models [62, 63, 64]. In its simplest form, an induction head copies from repeating tokens in the context to make predictions. For example, when presented with a token stream $[A][B]\dots[A]$, an induction head will output $[B]$ as the next token. Prior works train transformers on minimal Markov chain grammars, and find that, during pretraining, models learn to identify increasingly higher-order $k$-grams, with different attention heads specializing in copying, lookup, and aggregation [65, 66]. These works imply that pretraining enables models to learn conditional distributions, allowing them to represent sequence distributions seen in the context [67, 68, 69, 70].
@@
There is a clear parallel between context parroting and induction heads: both are essentially copy-and-paste operations, with context parroting involving the matching of not just one but multiple tokens. In fact, it is easy to imagine context parroting emerging naturally from combining multiple induction heads. This parallel can potentially explain the unreasonable effectiveness of applying language models trained on text to time-series tasks without fine-tuning or prompt engineering [27, 29, 30, 31, 61]. The induction heads formed from training on natural language happen to be also effective for predicting time series and can be easily repurposed to implement strategies such as context parroting.
3 Context parroting as an effective strategy for zero-shot forecasting
Overview of context parroting. In this section we motivate and introduce our baseline: context parroting. It was first identified as a strategy used by Chronos to predict chaotic systems, such as the partially observed Lorenz system shown in Fig. 1. It was also shown that one can predict the forecast accuracy of Chronos based on the context trajectory [42]: If there is a motif in the context that closely matches what happened immediately before the prediction starts, then Chronos would do well. This suggests that Chronos heavily utilizes context parroting to forecast chaotic systems. The parroting mechanism also explains Chronos’s sensitivity to initial conditions and sampling rate.
@@ 3 Context parroting as an effective strategy for zero-shot forecasting
Figure 1: Example of a foundation model forecasting chaotic dynamics with context parroting. Here, the foundation model (Chronos) was asked to predict the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points. Blue is the prediction and gray is the ground truth. Chronos produced an accurate prediction by simply looking for a motif in the context similar to the motif immediately preceding the prediction (highlighted in yellow) and copying the evolution following the matching motif (highlighted by pink boxes). We distill this context parroting strategy into Algorithm 1 and compare it against time-series foundation models (including Chronos itself).
@@
Context parroting is a simple forecasting strategy in which a model uses the last $D$ tokens of its context to query the remaining context. For whatever context sequence most closely matches the query, the following $H$ subsequent tokens are copied and used as the forecast. During the matching process, we exclude the last $D$ motifs to avoid parroting too close to where the prediction starts. Framed in terms of induction heads, the query lookup acts as a copy head, the nearest-neighbor match is a selector, and the exact repetion is the aggregation operation [66]. We describe this approach in Algorithm 1.
@@
Input: Context trajectory $x_{1:t}=\left\{x_{1},\ldots,x_{t}\right\}$, embedding dimension $D$ and forecast length $h$.
Output: Forecast trajectory $x_{t+1:t+h}=\left\{x_{t+1},\ldots,x_{t+h}\right\}$.
1:for all length-$D$ motif $s$: $x_{s-D+1:s}$ in the context $x_{1:t-D}$ do
2:     compute the distance $d_{s}$ between motif $s$ and the last motif $x_{t-D+1:t}$
3:Find the best-matching motif, $s_{opt}$, with the smallest distance
4:Set the first $t-s_{opt}$ predicted points to be $x_{t+1:2t-s_{opt}}=x_{s_{opt}+1:t}$ and repeat until the forecast length $h$ is reached
@@
In practice, the second kernel on $\mathbf{y}$ in Eq. 5 becomes a delta function, and so we write the conditional mean predictor
@@
Context parroting preserves attractor properties at long context lengths. In Appendix C.3, we derive the following proposition,
@@
Relationship to classical nonlinear forecasting. We show in Appendix C.2 that, in various limits, context parroting reduces to two classical algorithms from nonlinear dynamics: the simplex projection technique and the S-map algorithm [71, 72]. Both approaches have their foundations in Takens’ embedding theorem, which states that time-delayed observables derived from a nonlinear dynamical system will recover key geometric properties of the underlying attractor [73]. However, unlike context parroting, which looks for the best matching motif, simplex projection tries to identify multiple matching motifs in the context and computes a weighted average as its forecast. This makes simplex projection more sensitive to the choice of the embedding dimension (i.e., the length of the motif $D$), limiting the method to small embedding dimensions in practice [74].
@@ 4 Methods
Datasets. The dysts dataset provides a standardized benchmark of 135 low-dimensional chaotic systems, each defined by a set of ordinary differential equations between dimensionality three and six [75]. Every system is annotated with its largest Lyapunov exponent $\lambda$, an invariant characteristic of the underlying dynamics that quantifies the rate at which small perturbations grow over time. In chaotic systems, even minor errors rapidly compound over a characteristic timescale known as the Lyapunov time, defined as $\tau\equiv\lambda^{-1}$. As a result, highly chaotic systems (with small $\tau$) are especially challenging to forecast.
@@
Models. For time-series foundation models, we select Chronosbase (200M parameters), its variant Chronos-Boltbase (205M parameters), Time-MoElarge (200M parameters), and TimesFM-2.0 (500M parameters). All of these models are pretrained on massive amounts of real-world time series data (hundreds of billions of data points), which are often complemented by synthetic data to improve generalization. The three models encompass a wide array of design choices. Time-MoE and TimesFM employ decoder-only architectures, whereas Chronos adopts an encoder-decoder architecture. Chronos and Time-MoE use pointwise tokenization, whereas TimesFM chose patch tokenization. Time-MoE and TimesFM by default give point forecasts, whereas Chronos is designed to provide probabilistic forecasts with uncertainty quantification. Thus, for Chronos and Chronos-Bolt, we use the median prediction when evaluating forecast errors. An important parameter for these foundation models is the maximum context length $L_{\max}$. Chronos has the shortest allowed context window at $512$ data points, TimesFM-2.0 can handle $2048$ data points in the context, and Time-MoE has the largest capacity at $L_{\max}=4096$.
@@
Pipelines. To evaluate different models’ ability to zero-shot forecast dynamical systems, we generate a chaotic trajectory of length $10^{5}$ for each of the $135$ chaotic systems in dysts, with a granularity of $30$ data points per Lyapunov time. For a given context length $L$, we randomly pick a length-$L$ segment from the chaotic trajectory and provide it to the model as the context. The model’s task is to predict the next $300$ data points (equivalent to $10$ Lyapunov times) solely based on the context. We ask the model to make a univariate forecast on each dimension independently, which is then evaluated separately for each dimension. To obtain reliable statistics, we aggregate the results over all $135$ chaotic systems, all dimensions, and $20$ random initial conditions for each system.
@@
Metrics. In line with previous research [76, 77, 75, 78], we assess forecasting performance using two key metrics.
@@
where the sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ denotes the ground truth, and $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},\dots,\hat{\mathbf{x}}_{T}$ are the corresponding predictions made by the model.
@@
We use $\epsilon=30$, consistent with prior work [79, 78, 42].
In the appendix, we also show benchmark results using Mean Square Error (MSE) and Mean Absolute Error (MAE), two other metrics commonly used in the time series literature.

### v2

@@ Abstract
Recent time-series foundation models exhibit strong abilities to predict physical systems. These abilities include zero-shot forecasting, in which a model forecasts future states of a system given only a short trajectory as context, without knowledge of the underlying physics. Here, we show that foundation models often forecast through a simple parroting strategy, and when they are not parroting they exhibit some shared failure modes such as converging to the mean. As a result, a naive context parroting model that copies directly from the context scores higher than leading time-series foundation models on predicting a diverse range of dynamical systems, including low-dimensional chaos, turbulence, coupled oscillators, and electrocardiograms—and at a tiny fraction of the computational cost. We draw a parallel between context parroting and induction heads, which explains recent works showing that large language models can often be repurposed for time series forecasting. Our dynamical systems perspective also ties the scaling between forecast accuracy and context length to the fractal dimension of the underlying chaotic attractor, providing insight into previously observed in-context neural scaling laws. By revealing the performance gaps and failure modes of current time-series foundation models, context parroting can guide the design of future foundation models and help identify in-context learning strategies beyond parroting.
@@ 1 Introduction
A key test of generalization in scientific machine learning (SciML) is zero-shot forecasting: the ability to forecast future states of a new physical system based on a short context trajectory. Prior SciML approaches primarily focus on developing specialized forecasting models trained specifically on the system that needs to be predicted [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. However, the generality of these models is limited by the amount of system-specific data available, motivating the recent development of time-series foundation models [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], which are trained on vast amounts of observed and simulated time series from diverse domains, and which can subsequently perform zero-shot forecasts for any time series—including those generated by previously-unseen dynamical systems. Interestingly, it was recently found that, when available historical data is limited, time-series foundation models outperform classical deep learning models in forecasting chaotic dynamical systems [42].
What mechanisms do time-series foundation models use to make zero-shot forecasts, and why they are effective for dynamical systems not seen during pre-training? It was recently observed that one such foundation model, Chronos [35], often employs an extremely simple strategy when forecasting chaotic systems [42]. The strategy, context parroting, scans the context for nearly repeating motifs and copies the part of the context following the best-matching motif as its prediction (Fig.˜1). This can be viewed as a kind of “in-context nearest neighbor” algorithm, which is easy to implement during in-context computation [43]. How good is context parroting as a zero-shot forecasting strategy? By comparing it with existing foundation models, what can we learn about current models’ strengths and limitations?
Here, we compare context parroting with three state-of-the-art time-series foundation models: Chronos and Chronos Bolt [35], TimesFM [37], and Time-MoE [39] on the challenging task of forecasting chaotic systems. We find that context parroting outperforms all three foundation models in both zero-shot forecast accuracy and inference cost, especially for longer context windows. Our results suggest that current time-series foundation models do not fully utilize the information in the context data, and thus still have significant room for improvement when it comes to SciML tasks.
@@
    Introduce context parroting as a simple but effective baseline for zero-shot forecasting of dynamical systems, which can guide the design of more informative benchmarks that cannot be solved by simple repetitions and help identify forecasting strategies beyond parroting
@@
    Show that context parroting outperforms time-series foundation models in predicting chaotic systems and reveal common failure modes of the foundation models, which can guide the design of better models in the future
@@
    Explain the in-context neural scaling law between forecast accuracy and context length, linking the scaling coefficient to the fractal dimension of the underlying chaotic attractor
@@ 2 Related work
Foundation models for science. Foundation models have recently been introduced for many scientific machine-learning tasks [44], including partial differential equations [45, 46, 47, 48, 49, 50, 51], neuroscience [52, 53, 54], and weather forecasting [55, 56]. However, most of these foundation models remain a black box, and they have not yet provided interpretable strategies for forecasting diverse physical and dynamical processes. Here, we analyze context parroting as a simple mechanism used by time-series foundation models, noting its strengths and weaknesses as a zero-shot forecasting strategy. This strategy, and the insights gained here, can potentially be applied to other scientific tasks.
@@
In-context neural scaling laws. Neural scaling laws describe the relationship between the performance of a neural network and certain resources, such as model size, data size, or the amount of compute [57, 58, 59, 60]. Such scaling laws allow practitioners to predict the performance of yet-to-be-trained models based on the available resources and allocate them strategically to optimize compute-adjusted accuracy [61]. When applying LLMs to forecast dynamical systems, Liu et al. [62] recently observed an in-context neural scaling law, in which the test loss decreases with the context length following a power law. Here, we show that this in-context neural scaling law can be reproduced when using context parroting to predict dynamical systems, and the scaling coefficient can be linked to an invariant property of the underlying dynamic process (the fractal dimension of the chaotic attractor). This finding shows that neural scaling laws are intrinsically linked to invariant properties of the process generating the data, and the theory can potentially be generalized to other models and tasks (e.g., can we estimate the “fractal dimension” of a language from the neural scaling laws of LLMs?).
@@
In-context learning and induction heads. Induction heads are computational circuits that naturally emerge in simple transformers through training, and they have been hypothesized to underlie much of the in-context learning ability of foundation models [63, 64, 65, 66]. In its simplest form, an induction head copies repeating tokens in the context to make predictions. For example, when presented with a token stream $A][B]\dots[A]$, an induction head will output $[B]$ as the next token. Prior works train transformers on minimal Markov chain grammars, and find that, during pretraining, models learn to identify increasingly higher-order $k$-grams, with different attention heads specializing in copying, lookup, and aggregation [67, 68]. These works imply that pretraining enables models to learn conditional distributions, allowing them to represent sequence distributions seen in the context [69, 70, 71, 72].
@@
There is a clear parallel between context parroting and induction heads: both are essentially copy-and-paste operations, with context parroting involving the matching of not just one but multiple contiguous tokens. In fact, it is easy to imagine context parroting emerging naturally from combining multiple induction heads. This parallel can potentially explain the unreasonable effectiveness of applying language models trained on text to time-series tasks without fine-tuning or prompt engineering [27, 29, 30, 31, 62]. The induction heads formed from training on natural language happen to be also effective for predicting time series and can be easily repurposed to implement strategies such as context parroting.
3 Context parroting as a zero-shot forecasting strategy
Overview of context parroting. In this section we motivate and introduce our baseline: context parroting. It was inspired by recent observations that Chronos often predicts chaotic systems by copying directly from the context [42]. An example of Chronos using parroting to forecast a partially-observed Lorenz system is shown in Fig.˜1.
@@ 3 Context parroting as a zero-shot forecasting strategy
Figure 1: Example of a foundation model forecasting chaotic dynamics with context parroting. Here, the foundation model (Chronos) was asked to predict the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points. Blue is the prediction and gray is the ground truth. Chronos produced an accurate prediction by simply looking for a motif in the context similar to the motif immediately preceding the prediction (highlighted in yellow) and copying the evolution following the matching motif (highlighted by pink boxes). We distill this context parroting strategy into Algorithm˜1 and compare it against time-series foundation models (including Chronos itself).
@@
On a high level, context parroting uses the last $D$ tokens of the context to query the remaining context. For whatever context sequence that most closely matches the query, the subsequent tokens in the context are copied and used as the forecast. Because the length of the motif $D$ can be seen as the number of delayed states in a delay embedding from the lens of Takens’ embedding theorem [73], we also refer to $D$ as the embedding dimension and will use the terms embedding dimension and query length interchangeably. Interpreting $D$ as the embedding dimension is convenient because context parroting can be seen as a nearest neighbor algorithm in the $D$-dimensional delay-embedded space. During the matching process, we exclude the last $D$ motifs to avoid parroting too close to where the prediction starts. Framed in terms of induction heads, the query lookup acts as a copy head, the nearest-neighbor match is a selector, and the exact repetition is the aggregation operation [68]. We provide a pseudocode for context parroting in Algorithm˜1.
@@
Input: Context trajectory $x_{1:L}=\left\{x_{1},\ldots,x_{L}\right\}$, embedding dimension $D$ (i.e., the length of the motif to match), and forecast length $H$.
Output: Forecast trajectory $x_{L+1:L+H}=\left\{x_{L+1},\ldots,x_{L+H}\right\}$.
1:for all length-$D$ motif $s$: $x_{s-D+1:s}$ in the context $x_{1:L-D}$ do
2:  compute the Euclidean distance $d_{s}$ between motif $s$ and the last motif $x_{L-D+1:L}$
3:Find the best-matching motif, $s_{opt}$, with the smallest Euclidean distance
4:Set the first $L-s_{opt}$ predicted points to be $x_{L+1:2L-s_{opt}}=x_{s_{opt}+1:L}$ and repeat until the forecast length $H$ is reached
@@
We set the second kernel on $\mathbf{y}$ in Eq. 1 to a delta function, in order to output predictions that exactly match sequences from the context, rather than nearby sequences in a least-squares sense. We write the conditional mean predictor
@@
Context parroting preserves attractor properties at long context lengths. In Appendix F.3, we derive the following proposition,
@@
Relationship to classical nonlinear forecasting methods. We show in Appendix F.2 that, in various limits, context parroting is equivalent to two classical algorithms from nonlinear dynamics: the simplex projection technique and the S-map algorithm [74, 75]. Both approaches have their foundations in Takens’ embedding theorem, which states that time-delayed low-dimensional observables derived from a nonlinear dynamical system can recover key geometric properties of the underlying high-dimensional attractor [73]. However, unlike context parroting, which looks for the best matching motif, simplex projection tries to identify multiple matching motifs in the context and computes a weighted average as its forecast. This can potentially make simplex projection more sensitive to the choice of the embedding dimension $D$, limiting the method to small embedding dimensions in practice [76]. Other than simplex projection and S-map, there also exists other zero-shot forecasting strategies from nonlinear dynamics. For example, the Farmer-Sidorowich method [77] looks at multiple nearest neighbors in the context and builds a local linear model to make forecasts. An interesting future direction is to compare these methods from nonlinear dynamics with context parroting and time-series foundation models, which might inspire new zero-shot forecasting strategies.
@@ 4 Methods
Datasets. The dysts dataset provides a standardized benchmark of 135 low-dimensional chaotic systems, each defined by a set of ordinary differential equations between dimensionality three and six [78]. The chaotic systems are drawn from different published papers and span fields such as neuroscience, climate science, fluid dynamics, and astrophysics. Every system is annotated with its largest Lyapunov exponent $\lambda$, an invariant characteristic of the underlying dynamics that quantifies the rate at which small perturbations grow over time. In chaotic systems, even minor errors rapidly compound over a characteristic timescale known as the Lyapunov time, defined as $\tau\equiv\lambda^{-1}$. To normalize the difficulty of predicting different chaotic systems (so results from the $135$ systems can be meaningfully compared and combined), we generate trajectory data with a fixed sampling rate of $30$ points per Lyapunov time, and also measure the forecast performance in terms of Lyapunov times. To show the relevance of our findings to a broad class of SciML tasks, later we also go beyond low-dimensional chaotic systems and simulated data by benchmarking on real datasets from ECG measurements and electronic circuits.
@@
Models. For time-series foundation models, we select Chronosbase (200M parameters), its variant Chronos-Boltbase (205M parameters), Time-MoElarge (200M parameters), and TimesFM-2.0 (500M parameters). All of these models are pretrained on massive amounts of real-world time series data (hundreds of billions of data points), which are often complemented by synthetic data to improve generalization. The three models encompass a wide array of design choices. Time-MoE and TimesFM employ decoder-only architectures, whereas Chronos adopts an encoder-decoder architecture. Chronos and Time-MoE use pointwise tokenization, whereas TimesFM chose patch tokenization. Time-MoE and TimesFM by default give point forecasts, whereas Chronos is designed to provide probabilistic forecasts with uncertainty quantification. Thus, for Chronos and Chronos-Bolt, we use the median prediction when evaluating forecast errors. An important parameter for these foundation models is the maximum context length $L_{\max}$. Chronos has the shortest allowed context window at $512$ data points, TimesFM-2.0 can handle $2048$ data points, and Time-MoE has the largest capacity at $L_{\max}=4096$.
@@
Pipelines. To evaluate different models’ ability to zero-shot forecast dynamical systems, we generate a chaotic trajectory of length $10^{5}$ for each of the $135$ chaotic systems in dysts, with a granularity of $30$ data points per Lyapunov time. Each trajectory is normalized to have zero mean and unit standard deviation. For a given context length $L$, we randomly pick a length-$L$ segment from the chaotic trajectory and provide it to the model as the context. The model’s task is to predict the next $300$ data points (equivalent to $10$ Lyapunov times) solely based on the context. We ask the model to make a univariate forecast on each dimension independently, which is then evaluated separately for each dimension. To obtain reliable statistics, we aggregate the results over all $135$ chaotic systems, all dimensions, and $20$ random initial conditions for each system.
@@
Metrics. In line with previous research [79, 80, 78, 81], we assess forecasting performance using a diverse set of complementary metrics.
@@
where the sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ denotes the ground truth, and $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},\dots,\hat{\mathbf{x}}_{T}$ are the corresponding predictions made by the model. To provide some context to help interpret the sMAPE value, we note that predicting the mean of white noise would give you an sMAPE around $200$.
@@
We use $\epsilon=30$, consistent with prior work [82, 81, 42]. For chaotic systems, consistently achieving a VPT over one Lyapunov time is considered impressive [81, 42].
Other than sMAPE and VPT, we also show benchmark results using Mean Square Error (MSE) and Mean Absolute Error (MAE), two other metrics commonly used in the time series literature.
For chaotic dynamical systems, point forecasts will inevitably fail due to the exponential rate of error accumulation. It is thus equally important for a forecasting model to preserve the long-term statistical and geometric properties of the chaotic attractors, such as Lyapunov exponents and the attractor dimension. Here, we compare the structure of true and predicted attractors by calculating the KL Divergence between their distributions.
Kullback–Leibler Divergence between Attractors ($D_{\mathrm{stsp}}$).
$D_{\mathrm{stsp}}\equiv D_{\mathrm{KL}}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\frac{P(x)}{Q(x)},$
where $P$ and $Q$ represent the true and the predicted attractor, respectively. When estimating $D_{\mathrm{stsp}}$, we follow the methodology in [83, 84]. Specifically, we place Gaussian kernels at each point in the true and predicted trajectories and estimate the KL divergence between these Gaussian mixtures using a sampling-based approximation [85].
In the appendix, we also measure attractor reconstruction accuracy using Correlation Dimension. The correlation dimension non-parametrically estimates the fractal dimension from a time series, by calculating the scaling of the number of attractor points that fall within a given radius of each point.